OpenStudy (anonymous):

\[\lim_{n\rightarrow \infty}{^n\sqrt{2^{1+3n}}}\]

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty}\sqrt[n]{2^{1+3n}}\] \[=\lim_{n \rightarrow \infty}(2^{1+3n})^{1/n}\] \[=\lim_{n \rightarrow \infty}(2^{1/n+3})\] apply the limit \[2^{0+3}\] \[=2^{3}=16\]

OpenStudy (anonymous):

actually, \[2^3 = 8\] but you got my answer yay thanks so much